# The Mass Distribution of Stars Stripped in Binaries: The Effect of Metallicity

### v1

@@ 5.4 Can rapidly rotating stars populate the helium star desert?
It could be possible to look for these objects and test if they exist using our stripped star mass distribution. If more than ten helium stars more massive than are found in very metal-poor environments with constant star-formation rates of /yr (e.g., ), it is not expected from either binary envelope-stripping or wind-stripping. Such objects would thus provide excellent candidates for exploring chemically homogeneous evolution.
@@ 6 Summary and Conclusions
In this work, we explored the effect of metallicity on the formation of hot stripped stars through envelope-stripping in binary stars. To do this, we used a Monte Carlo binary population synthesis code created specifically for binary-stripped stars (see § 2.2) and based on detailed single and binary stellar evolution models (see § 2.1). We summarize our findings below:
@@
We estimate the number of hot stripped stars that should exist in a population as a function of mass (see § 3.2, Table 1, and Fig. 5). Independent of metallicity, we find that stripped stars per of constant star formation should be present and sufficiently massive to reach core collapse (, Tauris et al. 2015). Similarly, we find that stripped stars with mass should exist.
@@
Scaling the number estimates with appropriate star-formation rates, we estimate the number of hot stripped stars in the Milky Way, LMC, and SMC (see § 4). We find that there should be around 8,000 (60,000), 800 (6,000), and 200 (1,600) stripped stars in the Milky Way, LMC, and SMC, respectively, and with mass above 2.6 (1).
@@
We study the shape of the mass distribution for hot stripped stars (see § 3.1). At the low-mass end, the slope of the present-day mass function ( to ) is similar for all four metallicities that we consider, and it is also shallower than the slope of the adopted initial mass function (). This slope is sensitive to, for example, the assumed binary fraction and the adopted prescription for main sequence core overshoot. We discuss how, if measured from observations, this slope could potentially be used to help constrain these uncertain properties.
@@
We find that low-metallicity environments are expected to host many fewer massive hot stripped stars than high-metallicity environments. The reason is that, at low metallicity, massive stars are predicted to expand predominantly only after central helium burning has started, delaying binary interaction. The resulting lack of massive stripped stars can be seen in our computed mass function as a drop at the high-mass end of the distribution (see Fig. 4). For example, we estimate that only hot binary-stripped stars with should exist in the SMC (§ 4.3). Of these, we only expect to be sufficiently massive to exhibit a WR-type spectrum. Our results thus suggest that binary interaction is not efficient at forming WR stars at low metallicity.
@@
This lack of massive hot stripped stars at low metallicity can be described as a “helium star desert” — a triangular region of mass-metallicity space with and ranging from (see Fig. 6 and § 3.3), where our models predict that helium stars neither form efficiently through binary interaction nor through wind mass loss. If an unexpected amount ( per 1/yr at ) of helium stars are found in this region of mass-metallicity space, it is possible that they are rapidly-rotating stars that evolve chemically homogeneously (see § 5.4). However, it could also mean that the predicted late expansion at low metallicity does not, in fact, occur in nature, and that additional work is required to constrain and model the post-main sequence evolution of massive stars across metallicities.
@@
For example, the predicted lack of massive hot stripped stars at low metallicity could limit their contribution to the hard ionizing radiation observed in high- galaxies and their local analogs, as well as their contribution to cosmic reionization (cf. Götberg et al., 2020a). We made a rough estimate of the ionizing contribution from massive stripped stars (see § 5.2). We found that while they are non-negligible contributors of H-ionizing photons, stripped stars could potentially dominate the output of He+-ionizing photons from galaxies, even with the significant drop in their mass distribution. To better understand their roles, more careful spectral modeling and better understanding of wind driving from hot stripped stars are needed (such as Vink, 2017; Sander & Vink, 2020), as are observations that can confirm or refute theoretical predictions.
In addition, if it is difficult for massive stripped stars to form at low metallicity, this could also affect the efficiency of the conventionally-assumed isolated binary evolution pathway to the formation of compact object binaries. Here, we did not explore the impact of late expansion on compact object formation in depth, since it would require a larger exploration of detailed binary evolution models.
@@ Acknowledgements.
BHA thanks the Caltech Summer Undergraduate Research Fellowship (SURF) program and Peter Adams for supporting this project in memory of Alain Porter and Arthur R. Adams. BHA thanks Gwen Rudie for organizing the Carnegie Astrophysics Summer Student Internship (CASSI) program and all the staff at Carnegie Observatories who help to support this program. MRD acknowledges support from the NSERC through grant RGPIN-2019-06186, the Canada Research Chairs Program, and the Dunlap Institute at the University of Toronto. We thank Tomer Shenar and Selma de Mink for the interesting discussions that helped us improve the content of § 4. Computing resources used for this work were made possible by a grant from the Ahmanson Foundation. We thank the Observatories of the Carnegie Institution for Science for support, including Chris Burns for help with computations. BHA is supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-2234667. This work made use of MESA (version 8118, Paxton et al. 2011, 2013, 2015, 2018, 2019; Jermyn et al. 2023), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), and SciPy (Virtanen et al., 2020).References
@@
- Astropy Collaboration et al. (2022) Astropy Collaboration, Price-Whelan, A. M., Lim, P. L., et al. 2022, ApJ, 935, 167
- Astropy Collaboration et al. (2018) Astropy Collaboration, Price-Whelan, A. M., Sipőcz, B. M., et al. 2018, AJ, 156, 123
- Astropy Collaboration et al. (2013) Astropy Collaboration, Robitaille, T. P., Tollerud, E. J., et al. 2013, A&A, 558, A33
@@
- Deshmukh et al. (2024) Deshmukh, K., Sana, H., Mérand, A., et al. 2024, arXiv e-prints, arXiv:2409.15212
@@
- Laplace et al. (2024) Laplace, E., Schneider, F. R. N., & Podsiadlowski, P. 2024, arXiv e-prints, arXiv:2409.02058
@@
- Ramachandran et al. (2024) Ramachandran, V., Sander, A. A. C., Pauli, D., et al. 2024, arXiv e-prints, arXiv:2406.17678
@@
- Romagnolo et al. (2024) Romagnolo, A., Klencki, J., Vigna-Gomez, A., & Belczynski, K. 2024, arXiv e-prints, arXiv:2410.17315
@@
- Virtanen et al. (2020) Virtanen, P., Gommers, R., Oliphant, T. E., et al. 2020, Nature Methods, 17, 261
@@ Appendix A Stripped stars formed through different channels
As outlined in § 2.2 and visualized in Fig. 2, there are several evolutionary pathways that lead to stripped star formation through envelope-stripping in binaries. Here, we consider mass transfer initiated during the main-sequence evolution, mass transfer initiated during the Hertzsprung gap, and successful common envelope ejection that was initiated during the Hertzsprung gap evolution of the donor.
@@
The binary evolutionary models that we use as a base to model the population of stripped stars are created for Case B type mass transfer, initiated rather early during the Hertzsprung gap passage (§ 2.1, Götberg et al. 2018). This relies on the assumption that stripped stars created through different envelope-stripping mechanisms should look similar to those produced through Case B mass transfer (see however Yoon et al. 2017; Gilkis et al. 2019). While this is a simplification, it is likely more accurate than the assumption that all stripped stars can be represented by pure helium stars, as is the case in most binary population synthesis codes. Stripped stars created through common envelope ejection are thought to have been more deeply stripped than stripped stars produced through mass transfer (e.g., Ivanova 2011; Götberg et al. 2017; Vigna-Gómez et al. 2022). Stripped stars produced through Case A mass transfer are thought to have lower masses than those produced through Case B, since the helium core is not fully defined until the main-sequence is complete. Case A mass transfer therefore causes immediate recession of the central convective region, thus redefining the core (e.g., Yoon et al. 2010). While these predictions for the dependence of stripped star mass on the different envelope-stripping mechanisms have yet to be observationally constrained, we note that the lower masses predicted for Case A stripped stars would steepen the slope of the mass distribution, further enhancing the dearth of massive helium stars.

### v2

@@ 5.4 Rapidly rotating stars and the helium star desert
It could be possible to look for these objects and test if they exist using our stripped star mass distribution. If more than ten helium stars more massive than are found in very metal-poor environments with constant star formation rates of /yr (e.g., ), it is not expected from either binary envelope-stripping or wind-stripping. Such objects would thus provide excellent candidates for exploring chemically homogeneous evolution.
@@ 6 Summary and conclusions
In this work, we explored the effect of metallicity on the formation of hot stripped stars through envelope-stripping in binary stars. To do this, we used a Monte Carlo binary population synthesis code created specifically for binary-stripped stars (see Sect. 2.2) and based on detailed single and binary stellar evolution models (see Sect. 2.1). We summarize our findings below:
@@
We estimated the number of hot stripped stars that should exist in a population as a function of mass (see Sect. 3.2, Table 1, and Fig. 5). Independent of metallicity, we found that stripped stars per of constant star formation should be present and sufficiently massive to reach core collapse (, Tauris et al. 2015). Similarly, we found that stripped stars with mass should exist.
@@
Scaling the number estimates with appropriate star formation rates, we estimated the number of hot stripped stars in the Milky Way, LMC, and SMC (see Sect. 4). We found that there should be around 8,000 (60,000), 800 (6,000), and 200 (1,600) stripped stars in the Milky Way, LMC, and SMC, respectively, and with mass above 2.6 (1).
@@
We studied the shape of the mass distribution for hot stripped stars (see Sect. 3.1). At the low-mass end, the slope of the present-day mass function ( to ) is similar for all four metallicities that we considered, and it is also shallower than the slope of the adopted initial mass function (). This slope is sensitive to, for example, the assumed binary fraction and the adopted prescription for main sequence core overshoot. We discuss how, if measured from observations, this slope could potentially be used to help constrain these uncertain properties.
@@
We found that low-metallicity environments are expected to host many fewer massive hot stripped stars than high-metallicity environments. The reason is that, at low metallicity, massive stars are predicted to expand predominantly only after central helium burning has started, delaying binary interaction. The resulting lack of massive stripped stars can be seen in our computed mass function as a drop at the high-mass end of the distribution (see Fig. 4). For example, we estimated that only hot binary-stripped stars with should exist in the SMC (Sect. 4.3). Of these, we only expect to be sufficiently massive to exhibit a WR-type spectrum. Our results thus suggest that binary interaction is not efficient at forming WR stars at low metallicity.
@@
This lack of massive hot stripped stars at low metallicity can be described as a “helium star desert” — a triangular region of mass-metallicity space with and ranging from (see Fig. 6 and Sect. 3.3), where our models predict that helium stars neither form efficiently through binary interaction nor through wind mass loss. If an unexpected amount ( per 1/yr at ) of helium stars are found in this region of mass-metallicity space, it is possible that they are rapidly rotating stars that evolve chemically homogeneously (see Sect. 5.4). However, it could also mean that the predicted late expansion at low metallicity does not, in fact, occur in nature, and that additional work is required to constrain and model the post-main sequence evolution of massive stars across metallicities.
@@
For example, the predicted lack of massive hot stripped stars at low metallicity could limit their contribution to the hard ionizing radiation observed in high- galaxies and their local analogs, as well as their contribution to cosmic reionization (cf. Götberg et al., 2020a). We made a rough estimate of the ionizing contribution from massive stripped stars (see Sect. 5.2). We found that while they are non-negligible contributors of H-ionizing photons, stripped stars could potentially dominate the output of He+-ionizing photons from galaxies, even with the significant drop in their mass distribution. To better understand their roles, more careful spectral modeling and better understanding of wind driving from hot stripped stars are needed (such as Vink, 2017; Sander & Vink, 2020), as are observations that can confirm or refute theoretical predictions.
In addition, if it is difficult for massive stripped stars to form at low metallicity, this could also affect the efficiency of the conventionally assumed isolated binary evolution pathway to the formation of compact object binaries. Here, we did not explore the impact of late expansion on compact object formation in depth, since it would require a larger exploration of detailed binary evolution models.
@@ Acknowledgements.
We thank the anonymous referee for providing a constructive report. We thank Tomer Shenar and Selma de Mink for the interesting discussions that helped us improve the content of Sect. 4. Thank you to Jorick Vink and Andreas Sander for helpful discussions about wind driving. BHA thanks the Caltech Summer Undergraduate Research Fellowship (SURF) program and Peter Adams for supporting this project in memory of Alain Porter and Arthur R. Adams. BHA thanks Gwen Rudie for organizing the Carnegie Astrophysics Summer Student Internship (CASSI) program and all the staff at Carnegie Observatories who help to support this program. BHA also thanks Laura Jaliff, Sal Wanying Fu, Ivanna Escala, Johanna Teske, Tony Piro, Brian Lorenz, and Peter Senchyna for their mentorship during this project. Computing resources used for this work were made possible by a grant from the Ahmanson Foundation. We thank the Observatories of the Carnegie Institution for Science for support, including Chris Burns for help with computations. This work used computing resources provided by Northwestern University and the Center for Interdisciplinary Exploration and Research in Astrophysics (CIERA). This research was supported in part through the computational resources and staff contributions provided for the Quest high performance computing facility at Northwestern University which is jointly supported by the Office of the Provost, the Office for Research, and Northwestern University Information Technology. MRD acknowledges support from the NSERC through grant RGPIN-2019-06186, the Canada Research Chairs Program, and the Dunlap Institute at the University of Toronto. BHA is supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-2234667.References
@@
- Bavera et al. (2021) Bavera, S. S., Fragos, T., Zevin, M., et al. 2021, A&A, 647, A153
@@
- Deshmukh et al. (2024) Deshmukh, K., Sana, H., Mérand, A., et al. 2024, A&A, 692, A109
@@
- Inayoshi et al. (2017) Inayoshi, K., Hirai, R., Kinugawa, T., & Hotokezaka, K. 2017, MNRAS, 468, 5020
@@
- Laplace et al. (2020) Laplace, E., Götberg, Y., de Mink, S. E., Justham, S., & Farmer, R. 2020, A&A, 637, A6
@@
- Laplace et al. (2025) Laplace, E., Schneider, F. R. N., & Podsiadlowski, P. 2025, A&A, 695, A71
@@
- Ramachandran et al. (2024) Ramachandran, V., Sander, A. A. C., Pauli, D., et al. 2024, A&A, 692, A90
@@
- Romagnolo et al. (2025) Romagnolo, A., Klencki, J., Vigna-Gómez, A., & Belczynski, K. 2025, A&A, 693, A137
@@ Appendix A Stripped stars formed through different channels
As outlined in Sect. 2.2 and visualized in Fig. 2, there are several evolutionary pathways that lead to stripped star formation through envelope-stripping in binaries. Here, we consider mass transfer initiated during the main-sequence evolution, mass transfer initiated during the Hertzsprung gap, and successful common envelope ejection that was initiated during the Hertzsprung gap evolution of the donor.
@@
The binary evolutionary models that we use as a base to model the population of stripped stars are created for Case B type mass transfer, initiated rather early during the Hertzsprung gap passage (Sect. 2.1, Götberg et al. 2018). This relies on the assumption that stripped stars created through different envelope-stripping mechanisms should look similar to those produced through Case B mass transfer (see however Yoon et al. 2017; Gilkis et al. 2019). While this is a simplification, it is likely more accurate than the assumption that all stripped stars can be represented by pure helium stars, as is the case in most binary population synthesis codes. Stripped stars created through common envelope ejection are thought to have been more deeply stripped than stripped stars produced through mass transfer (e.g., Ivanova 2011; Götberg et al. 2017; Vigna-Gómez et al. 2022). Stripped stars produced through Case A mass transfer are thought to have lower masses than those produced through Case B, since the helium core is not fully defined until the main-sequence is complete. Case A mass transfer therefore causes immediate recession of the central convective region, thus redefining the core (e.g., Yoon et al. 2010). While these predictions for the dependence of stripped star mass on the different envelope-stripping mechanisms have yet to be observationally constrained, we note that the lower masses predicted for Case A stripped stars would steepen the slope of the mass distribution, further enhancing the dearth of massive helium stars.
Appendix B Properties derived from MESA stellar evolution models
[…]
| 2.00 | 0.22 | 0.35 | – | |
| 2.21 | 0.26 | 0.38 | ||
| 2.44 | 0.29 | 0.44 | – | |
| 2.70 | 0.33 | 0.51 | ||
| 2.99 | 0.38 | 0.58 | – | |
| 3.30 | 0.44 | 0.67 | ||
| 3.65 | 0.50 | 0.74 | ||
| 4.04 | 0.57 | 0.85 | ||
| 4.46 | 0.65 | 0.97 | ||
| 4.93 | 0.76 | 1.12 | ||
| 5.45 | 0.88 | 1.28 | ||
| 6.03 | 1.01 | 1.43 | ||
| 6.66 | 1.18 | 1.64 | ||
| 7.37 | 1.38 | 1.89 | ||
| 8.15 | 1.61 | 2.18 | ||
| 9.00 | 1.88 | 2.50 | ||
| 9.95 | 2.20 | 2.88 | ||
| 11.00 | 2.59 | 3.34 | ||
| 12.17 | 3.04 | 3.86 | ||
| 13.45 | 3.58 | 4.46 | ||
| 14.87 | 4.22 | 5.14 | ||
| 16.44 | 4.93 | 5.90 | ||
| 18.17 | 5.73 | 6.74 | ||
| 20.09 | 6.68 | – | – | |
| 22.21 | 7.73 | – | – | |
| 24.55 | 8.91 | – | – | |
| 27.13 | 10.20 | – | – | |
| 30.00 | 11.71 | – | – | |
| 33.16 | 13.37 | – | – | |
| 36.65 | 15.21 | – | – | |
| 40.52 | 17.22 | – | – | |
| 44.79 | 19.51 | – | – | |
| 49.51 | 22.08 | – | – | |
| 54.73 | 24.93 | – | – | |
| 60.50 | 28.00 | – | – | |
| 66.89 | 31.66 | – | – | |
| 73.94 | 35.64 | – | – | |
| 81.74 | 40.02 | – | – | |
| 90.38 | 44.64 | – | – | |
| 99.91 | 47.56 | – | – |
| (yr) | (yr) | |||
| 2.00 | 0.24 | 0.37 | – | |
| 2.21 | 0.27 | 0.42 | – | |
| 2.44 | 0.31 | 0.48 | – | |
| 2.70 | 0.35 | 0.54 | – | |
| 2.99 | 0.40 | 0.62 | ||
| 3.30 | 0.46 | 0.71 | ||
| 3.65 | 0.52 | 0.81 | ||
| 4.04 | 0.60 | 0.92 | ||
| 4.46 | 0.68 | 1.06 | ||
| 4.93 | 0.79 | 1.21 | ||
| 5.45 | 0.91 | 1.38 | ||
| 6.03 | 1.05 | 1.55 | ||
| 6.66 | 1.21 | 1.77 | ||
| 7.37 | 1.42 | 2.04 | ||
| 8.15 | 1.65 | 2.34 | ||
| 9.00 | 1.93 | 2.70 | ||
| 9.95 | 2.27 | 3.11 | ||
| 11.00 | 2.65 | 3.59 | ||
| 12.17 | 3.11 | 4.14 | ||
| 13.45 | 3.64 | 4.77 | ||
| 14.87 | 4.26 | 5.49 | ||
| 16.44 | 4.99 | 6.30 | ||
| 18.17 | 5.80 | 7.17 | ||
| 20.09 | 6.72 | – | – | |
| 22.21 | 7.80 | – | – | |
| 24.55 | 9.00 | – | – | |
| 27.14 | 10.39 | – | – | |
| 30.00 | 11.93 | – | – | |
| 33.16 | 13.64 | – | – | |
| 36.66 | 15.57 | – | – | |
| 40.53 | 17.75 | – | – | |
| 44.80 | 20.15 | – | – | |
| 49.53 | 22.84 | – | – | |
| 54.75 | 25.87 | – | – | |
| 60.52 | 29.19 | – | – | |
| 66.91 | 33.02 | – | – | |
| 73.97 | 37.23 | – | – | |
| 81.78 | 41.87 | – | – | |
| 90.41 | 47.06 | – | – | |
| 99.96 | 52.90 | – | – |
| (yr) | (yr) | |||
| 2.00 | 0.25 | 0.39 | – | |
| 2.21 | 0.29 | 0.45 | – | |
| 2.44 | 0.33 | 0.51 | – | |
| 2.70 | 0.37 | 0.59 | ||
| 2.99 | 0.42 | 0.67 | ||
| 3.30 | 0.48 | 0.76 | ||
| 3.65 | 0.55 | 0.87 | ||
| 4.04 | 0.62 | 1.01 | ||
| 4.46 | 0.72 | 1.15 | ||
| 4.93 | 0.82 | 1.30 | ||
| 5.45 | 0.94 | 1.50 | ||
| 6.03 | 1.09 | 1.68 | ||
| 6.66 | 1.26 | 1.93 | ||
| 7.37 | 1.47 | 2.22 | ||
| 8.15 | 1.71 | 2.54 | ||
| 9.00 | 1.98 | 2.92 | ||
| 9.95 | 2.32 | 3.37 | ||
| 11.01 | 2.71 | 3.88 | ||
| 12.17 | 3.18 | 4.47 | ||
| 13.45 | 3.70 | 5.15 | ||
| 14.87 | 4.33 | 5.91 | ||
| 16.44 | 5.05 | 6.63 | ||
| 18.17 | 5.85 | 7.26 | ||
| 20.09 | 6.80 | – | – | |
| 22.21 | 7.88 | – | – | |
| 24.55 | 9.12 | – | – | |
| 27.14 | 10.50 | – | – | |
| 30.00 | 12.08 | – | – | |
| 33.17 | 13.84 | – | – | |
| 36.66 | 15.84 | – | – | |
| 40.53 | 18.12 | – | – | |
| 44.81 | 20.62 | – | – | |
| 49.53 | 23.42 | – | – | |
| 54.76 | 26.55 | – | – | |
| 60.53 | 30.06 | – | – | |
| 66.92 | 33.96 | – | – | |
| 73.98 | 38.29 | – | – | |
| 81.79 | 43.13 | – | – | |
| 90.43 | 48.54 | – | – | |
| 99.97 | 54.54 | – | – |
| (yr) | (yr) | |||
| 2.00 | 0.26 | 0.42 | – | |
| 2.21 | 0.30 | 0.49 | – | |
| 2.44 | 0.34 | 0.57 | ||
| 2.70 | 0.39 | 0.66 | ||
| 2.99 | 0.44 | 0.74 | ||
| 3.30 | 0.50 | 0.87 | – | |
| 3.65 | 0.57 | 1.00 | ||
| 4.04 | 0.65 | 1.15 | ||
| 4.46 | 0.74 | 1.30 | ||
| 4.93 | 0.85 | 1.49 | ||
| 5.45 | 0.98 | 1.70 | ||
| 6.03 | 1.12 | 1.92 | ||
| 6.66 | 1.29 | 2.20 | ||
| 7.37 | 1.49 | 2.53 | ||
| 8.15 | 1.74 | 2.90 | ||
| 9.00 | 2.02 | 3.33 | ||
| 9.95 | 2.35 | 3.81 | ||
| 11.01 | 2.75 | 4.38 | ||
| 12.17 | 3.20 | 5.04 | ||
| 13.45 | 3.75 | 5.73 | ||
| 14.87 | 4.36 | 6.36 | – | |
| 16.44 | 5.08 | 7.20 | ||
| 18.17 | 5.90 | 7.92 | ||
| 20.09 | 6.86 | – | – | |
| 22.21 | 7.95 | – | – | |
| 24.55 | 9.19 | – | – | |
| 27.14 | 10.61 | – | – | |
| 30.00 | 12.20 | – | – | |
| 33.17 | 13.98 | – | – | |
| 36.67 | 16.00 | – | – | |
| 40.54 | 18.28 | – | – | |
| 44.81 | 20.84 | – | – | |
| 49.54 | 23.76 | – | – | |
| 54.77 | 26.96 | – | – | |
| 60.55 | 30.51 | – | – | |
| 66.94 | 34.50 | – | – | |
| 74.01 | 39.02 | – | – | |
| 81.82 | 43.90 | – | – | |
| 90.45 | 49.40 | – | – | |
| 99.99 | 55.46 | – | – |